Calculate the Kb of the conjugate base of a weak acid with a Ka = 5.28 × 10−5.

- report answer in three significant figures  
- answer should be written in scientific notation    
(ex. 4.67E-7 or 4.67E7 be sure to use a CAPITAL E not lower case 'e')
Answer #1

with Ka = 5.28x10-5 1.0x10-14 [ Ka ] 1.0x10-14 [ 5.28x10-5 Kb = 1.89x10-10 Kb = 1.89E-10 Kb =